OpenStudy (anonymous):

Two sequences have the same common difference. How many terms could the sequences have in common?

OpenStudy (cj49):

let the two equation be P1 and P2 with last terms as 100 and 90 repectively. Let the common difference be 5 P3. the last term of P3 cannot be greater than 90 let the 1st common no. be 11 \[11+ (n-1)5 \le 90\] \[(n-1)5\le90-11=81\] \[(n-1)\le80/5=17\] n=18

OpenStudy (anonymous):

Where did the numbers come from?

OpenStudy (cj49):

it was just an example

OpenStudy (ankit042):

as common diff is same so let us assume it is d starting term for first seq x starting term for second seq y Now there can be three case x>y y<x or x=y

OpenStudy (ankit042):

I think you can answer now

OpenStudy (anonymous):

Thank you c:

OpenStudy (ankit042):

np! what's answer did you get?

OpenStudy (anonymous):

I'm supposing that the answer can be infinitive since the sequence is

OpenStudy (anonymous):

@ankit042

OpenStudy (cj49):

the answer depends on the sequence

OpenStudy (ankit042):

Yes I agree in case x=y we will have all terms same but for other cases no terms will be same!

OpenStudy (anonymous):

So in another sequence will it not have similar terms?

OpenStudy (cj49):

a=1,3,5,7,8,10.......100 terms b=7,9,12......90 terms the number of common terms would change

OpenStudy (ankit042):

@jherrera98 sorry my bad in other cases also you can have same terms sorry I got confused

OpenStudy (anonymous):

Ohh ok c: thank you ! @cj49 @ankit042

OpenStudy (cj49):

np.